OpenStudy (anonymous):

HELP!!! sinx - 2 = cosx - 2. Solve for x.

OpenStudy (anonymous):

sin x = cos x = sin (Pi/2 - x) sin x = sin (Pi/2 - x) x = Pi/2 - x -> 2x = Pi/2 -> x = ...

OpenStudy (anonymous):

Refer to the attachment from Wolffram Alpha.

OpenStudy (anonymous):

sinx=cosx sinx/cosx=tanx=1 x=pi/4

OpenStudy (anonymous):

those value x which give 1 as output like pi/4 5pi/4 9pi/4